### v1

# Cover times with stochastic resetting

Cover times with stochastic resetting
Samantha Linn  and  Sean D. Lawley
Cover times quantify the speed of exhaustive search. In this work, we compute exactly the mean cover time associated with a one-dimensional Brownian search under exponentially distributed resetting. We also approximate the moments of cover times of a wide range of stochastic search processes in $d$-dimensional continuous space and on an arbitrary discrete network under frequent stochastic resetting. These results hold for a large class of resetting time distributions and search processes including diffusion and Markov jump processes.
SL: Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (slinn.math@gmail.com). SL was supported by the National Science Foundation (Grant No. 2139322)
SDL: Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (lawley@math.utah.edu). SDL was supported by the National Science Foundation (CAREER DMS-1944574 and DMS-2325258).
1. Introduction
An exhaustive search is one in which an entire state space is explored by a searcher. Such search processes are ubiquitous in technology and the natural world, ranging from computer search algorithms to animal foraging or pathogen hunting. In some cases the state space of interest, or target region, may be a proper subset of the searcher domain but it may also be the entire domain. Whatever the target region, we call the time taken to complete such a search its cover time [1].
To be precise, denote the position of a stochastic process in a domain $M\subseteq\mathbb{R}^{d}$ by $X=\{X(t)\}_{t\geq 0}$. If this stochastic ‘searcher’ has detection radius $R>0$, then the region it explores by time $t\geq 0$ is
$S_{0}(t):=\cup_{s=0}^{t}B(X(s),R)\subseteq M,$
where $B(X,R)$ is the ball of radius $R>0$ centered at $X\in M$. Hence, the cover time of a given target region $U_{T}\subseteq M$ is
$T_{0}:=\inf\{t>0:U_{T}\subseteq S_{0}(t)\}.$
The bulk of previous work on cover times concerns the case that the searcher detection radius vanishes compared to the size of the domain. In continuous space, it has been shown that cover times of Brownian motion diverge as the detection radius goes to zero [2]:
$T_{0}\sim t_{d}\quad\text{almost surely as }R\to 0,$
where $t_{d}$ depends on the searcher diffusivity, $D>0$, the domain volume, $|M|$, and the spatial dimension, $d>0$,
$t_{d}=\begin{cases}\frac{1}{\pi}\frac{|M|}{D}(\ln(1/R))^{2}\quad&\text{if }d=2,\\ \frac{d\Gamma(d/2)R^{2-d}}{2(d-2)\pi^{d/2}}\frac{|M|}{D}\ln(1/R)\quad&\text{if}d\geq 3.\end{cases}$
Throughout, $f\sim g$ denotes $f/g\to 1$. In discrete space, it was shown that cover times of random walks follow a Gumbel distribution and diverge in the limit of a large target region [3].
More recently, others have considered cover times of multiple simultaneous searchers of fixed detection radius in the many searcher limit. In continuous space, the $m$th moment of the cover time of $U_{T}\subseteq M$ for $N\gg 1$ Brownian searchers, $T(N)\geq 0$, satisfies
$\mathbb{E}((T(N))^{m})\sim\Big{(}\frac{L^{2}}{4D\ln N}\Big{)}^{m}\quad\text{as }N\to\infty,$
where $D>0$ denotes searcher diffusivity and $L>0$ denotes the distance the searcher must travel to reach the furthest part of the target region [4]. Moreover, in discrete space, these moments are given by
$\mathbb{E}((T(N))^{m})\sim\frac{\kappa_{m}}{N^{m/d}}\quad\text{as}N\to\infty,$
where $d\geq 1$ is the fewest steps the searcher must take to reach the furthest part of the target region and $\kappa_{m}$ depends on the jump rates along this path [5]. An important observation in the $N\gg 1$ searcher case is that the cover times can be closely approximated with merely some information about the geodesic paths from the initial position of the searcher to the furthest part(s) of the target region. The same can be said of the work herein.
In this paper we consider the cover time of a single stochastic searcher that undergoes stochastic resetting to its initial position. In continuous space the searcher has a fixed detection radius and in discrete space the searcher detects only the node it occupies. To our knowledge, the only other investigation of cover times under stochastic resetting is the work of Colombani et al [6] who studied how resetting can be used to reduce mean cover times (MCTs) of a self-avoiding random walker on a discrete network.
Not only is stochastic resetting a useful formalism to model physical phenomena [7, 8, 9], it can also expedite certain search processes [10, 11, 12]. For instance, consider the mean escape time of a Brownian searcher with diffusivity $D>0$ on the half line. It is well-known that without stochastic resetting the mean first passage time to the origin is infinite. That is, denoting its position at time $t\geq 0$ by $\{X(t)\}_{t\geq 0}$ and supposing $X(0)=x_{0}>0$, we have that
$\mathbb{E}[\text{inf}\{t>0:X(t)\leq 0\}]=\infty.$
However, if we now impose that this searcher instantaneously resets to its initial position according to an exponential random variable with rate $r>0$, the mean first passage time to the origin is finite,
$\mathbb{E}[\text{inf}\{t>0:X(t)\leq 0\}]=\frac{1}{r}(e^{x_{0}\sqrt{r/D}}-1)<\infty,$
and attains a minimum for a strictly positive resetting rate [13].
Similar to mean first passage times, we show that stochastic resetting can be used to reduce MCTs in continuous space. Moreover, with a well-defined notion of resetting rate, $r>0$, we show that the frequent resetting asymptotics of the moments of cover times in discrete and continuous space depend only on the geodesic path(s) from the initial (resetting) position of the searcher to the furthest part(s) of the target region. Our results hold for any fixed detection radius. While these results are exact only in the frequent resetting limit, we can compute exactly the mean cover time with any resetting rate $r>0$ in the special case of a one-dimensional Brownian searcher.
All these results are made precise in the rest of the paper, which is organized as follows. In section 2, we consider the one-dimensional Brownian cover time problem with resetting. In section 3, we consider the $d$-dimensional continuous space cover time problem with resetting. In section 4, we consider the cover time problem with resetting on an arbitrary discrete network. We conclude by discussing the scope of the results and proposing directions of future work.
2. Exact mean cover times for a resetting Brownian searcher in $\mathbb{R}$
Consider a Brownian searcher with diffusivity $D>0$ who resets according to exponential random variable with rate $r>0$ to its initial position. We denote its position at time $t\geq 0$ by $X_{r}(t)$ and denote its first hitting time to a set $\Omega$ with initial position $x_{0}$ and resetting position $x_{r}$ by $\tau_{r}(\Omega|x_{0},x_{r})$. If $x_{0}=x_{r}=0$, we abbreviate $\tau_{r}(\Omega)\equiv\tau_{r}(\Omega|0,0)$.
In this section, we determine the mean cover time of the interval $[-a,b]$ with $a>0$ and $b>0$. We consider two cases: In the first case the searcher is unconstrained on the real line and in the second case the searcher is confined to the interval $[-a,b]$ with reflecting boundary conditions. In section 3, we ascribe a detection radius to the searcher and study this system in higher dimensions and more general domains.
2.1. Unconstrained search
Consider first that the searcher is unconstrained on $\mathbb{R}$. It follows from the strong Markov property that the cover time of the interval $[-a,b]$ can be described by the sum of the exit time from $(-a,b)$ and the exit time from the half line with initial position $X_{r}(0)=a+b$ and resetting position either $a>0$ or $b>0$ depending on which interval boundary was first contacted. The mean interval exit time is known [14],
(1) $\mathbb{E}(\tau_{r}(\{-a,b\}))=\frac{1}{r}\left[\frac{\sinh((a+b)\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)}-1\right]$
where $\alpha:=\sqrt{r/D}$. The mean half line exit time with initial position $x_{0}>0$ and resetting position $x_{r}>0$ is also known [13],
(2) $\mathbb{E}(\tau_{r}(\{0\}|x_{0},x_{r})):=\mathbb{E}(\inf\{t>0:X_{r}(t)\not>0\})=\frac{1-e^{-\alpha x_{0}}}{re^{-\alpha x_{r}}}.$
Furthermore, the hitting probabilities of $x=-a$ and $x=b$ are given by [14]
(3) $\epsilon_{a}$ $:=\mathbb{P}(\tau_{r}(\{-a\})<\tau_{r}(\{b\}))=\frac{\sinh(b\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)},$
(4) $\epsilon_{b}$ $:=1-\mathbb{P}(\tau_{r}(\{-a\})<\tau_{r}(\{b\}))=\frac{\sinh(a\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)}.$
Denote the cover time by $T_{r}>0$. Combining equations 1, 3, 4 and 2 yields the MCT of the interval $[-a,b]$ by a resetting Brownian searcher,
(5) $\begin{split}\mathbb{E}(T_{r})&=\mathbb{E}(\tau_{r}(\{-a,b\}))+\epsilon_{a}\mathbb{E}(\tau_{r}(\{0\}|a+b,b))+\epsilon_{b}\mathbb{E}(\tau_{r}(\{0\}|a+b,a))\\ &=\frac{e^{(a+b)\alpha}+e^{2a\alpha}+e^{2b\alpha}-e^{a\alpha}-e^{b\alpha}-1}{r\left(e^{a\alpha}+e^{b\alpha}\right)}.\end{split}$
Notice in Figure 1 that for finite resetting rate $r$ the mean cover time is finite. This is in stark contrast to the analogous cover time problem without resetting; it is well-known that the mean first passage time to a point not equal to $X(0)$ is infinite and, hence, so too is the mean cover time of an interval of nonzero length. Observe further that the MCT of the interval $[-a,b]$ by a resetting Brownian motion on the (unconstrained) real line achieves a unique minimum with a positive resetting rate. This ‘optimal’ resetting rate, which is found by computing the unique zero of the $r$-derivative of (5), does not have an explicit analytical solution, but notice in Figure 1 how closely the frequent resetting result approximates the exact result.
Taking $r\to\infty$ in (5) reveals that the MCT depends on the relative sizes of $a>0$ and $b>0$. For instance, if $a<b$, one can verify that the large $r$ behavior of (5) behaves like
(6) $\mathbb{E}(T_{r})\sim\frac{1}{rp}\quad\text{as }r\to\infty$
where $p:=\text{exp}(-b\alpha)$ is the probability of hitting the further boundary point. In this case it is straightforward to compute the minimum MCT of the asymptotic result in (6) of $b^{2}e^{2}/4D$ at $r=4D/b^{2}$, which converges to the exact result exponentially fast in the square root of the resetting rate. Hence, the asymptotic result in (6) closely approximates the true optimum even for modest resetting rates. We find similar large $r$ behavior in much more general settings in section 3.
We briefly address the MCT of a resetting Brownian motion on a symmetric interval, i.e. the case $|a|=b$. Repeating the analysis above yields the same asymptotic result as in (6) with a prefactor of $3/2$. The same phenomenon can be observed in the case of a constrained search, which we consider in subsection 2.2 below. We give an intuitive explanation for this prefactor in the discussion.
2.2. Constrained search
Consider the same process but on the closed interval $[-a,b]$ with reflecting boundary conditions. The cover time of the interval can now be described by the sum of the hitting time of $x\in\{-a,b\}$ and the hitting time of $x=0$ on the closed interval $[0,a+b]$ with initial position $X_{r}(0)=a+b$ and resetting position either $a>0$ or $b>0$ depending on which interval boundary was first hit. The mean of the first term was computed in the previous section. Below we determine the mean time of the second term.
Suppose the searcher hits $x=b$ before $x=-a$. Denote the survival probability of this process with resetting rate $r>0$ by $Q_{r}(x,t)$,
$Q_{r}(x,t):=\mathbb{P}(\tau_{r}(\{-a\})>t|X_{r}(0)=x).$
Figure 1. MCTs of the interval $[-1,2]$ for a resetting Brownian motion with diffusivity $D=1$ and $L=2$ denoting the furthest target point from the searcher initial and resetting position. Resetting is exponentially distributed with rate $r>0$. See sections 2 and 3 for details.
The backward Fokker-Planck equation of the survival probability without resetting is given by
(7) $\frac{\partial Q_{0}(x,t)}{\partial t}=D\frac{\partial^{2}Q_{0}(x,t)}{\partial x^{2}}$
with boundary conditions $Q_{0}(-a,t)=0$ and $\frac{\partial Q_{0}(x,t)}{\partial x}\big{|}_{x=b}=0$. Denote the Laplace transform in time of $Q_{r}(x,t)$ by $q_{r}(x,s)$. The Laplace transform of (7) satisfies the ordinary differential equation
$sq_{0}(x,s)-1=Dq^{\prime\prime}_{0}(x,s)$
with boundary conditions $q_{0}(-a,s)=0$ and $\frac{\partial q_{0}(x,s)}{\partial x}\big{|}_{x=b}=0$. We substitute the solution of this boundary value problem into a known expression of $q_{r}(x,s)$ in terms of $q_{0}(x,s)$ (which can be derived by way of a renewal equation, see [13]),
(8) $q_{r}(x_{0},s)=\frac{q_{0}(x_{0},r+s)}{1-rq_{0}(x_{r},r+s)},$
and finally evaluate (8) at $s=0$ with $x_{0}=b$ and $x_{r}=0$ to find that
(9) $\mathbb{E}(\tau_{r}(\{-a\}|b,0))=\frac{(e^{(a+b)\alpha}-1)^{2}}{re^{a\alpha}(e^{2b\alpha}+1)}.$
By the analogous argument wherein the searcher hits $x=-a$ before $x=b$, we find
(10) $\mathbb{E}(\tau_{r}(\{b\}|-a,0))=\frac{(e^{(a+b)\alpha}-1)^{2}}{re^{-b\alpha}(e^{2a\alpha}+1)}.$
Combining equations 9 and 10 with the mean time of the first term and the relevant hitting probabilities computed in the previous system yields the MCT of the closed interval $[-a,b]$ by a resetting Brownian searcher,
(11) $\begin{split}\mathbb{E}(T_{r})=\frac{1}{r}&\left[\frac{\sinh((a+b)\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)}-1\right]\dots\\ &\qquad+\frac{\epsilon_{a}}{r}\left[\frac{(e^{(a+b)\alpha}-1)^{2}}{e^{b\alpha}(e^{2a\alpha}+1)}\right]+\frac{\epsilon_{b}}{r}\left[\frac{(e^{(a+b)\alpha}-1)^{2}}{e^{a\alpha}(e^{2b\alpha}+1)}\right].\end{split}$
As in the unconstrained case, taking $r\to\infty$ reveals that the MCT depends on the relative sizes of $a>0$ and $b>0$. If $a<b$, the large $r$ behavior of (11) behaves like
$\mathbb{E}(T_{r})\sim\frac{1}{rp}\quad\text{as }r\to\infty$
where $p:=\text{exp}(-b\alpha)$ is again the probability of hitting the further boundary point before a resetting event. We illustrate this behavior in Figure 1. Similar to the unconstrained case, the asymptotic result here closely approximates the true optimum even for modest resetting rates.
3. Cover times for a frequently resetting searcher in $\mathbb{R}^{d}$
3.1. Definitions and main result
Now consider a stochastic searcher in the domain $M\subseteq\mathbb{R}^{d}$ with fixed detection radius $R>0$ which randomly resets to its initial position. We are most interested in the case of Brownian search, which we later show satisfies the minimal assumptions we impose on the search process. Denoting searcher position at time $t\geq 0$ by $X_{r}(t)$, the region detected by the searcher at time $t\geq 0$ given a metric $L:M\times M\to[0,\infty)$ is the closed ball of radius $R>0$ centered at $X_{r}(t)\in M$,
$B(X_{r}(t),R):=\{y\in M:L(X_{r}(t),y)\leq R\}.$
For Brownian search, $L$ is merely the standard Euclidean distance. We assume the initial and resetting position of the searcher is $X_{r}(0)=0$ and denote by $U_{0}$ the region it initially detects,
$U_{0}:=B(X_{r}(0),R)=B(0,R).$
We consider resetting processes, $\sigma>0$, that belong to the following family of random variables: Let $Y>0$ be a strictly positive random variable that does not depend on $r>0$ and has unit mean,
$\mathbb{E}[Y]=1.$
Moreover, suppose $Y>0$ has a finite moment generating function in a neighborhood of the origin,
$\mathbb{E}[e^{zY}]<\infty\quad\text{for all }z\in[-\delta,\delta]$
for some $\delta>0$. Define
(12) $\sigma:=Y/r.$
where we call $r>0$ the resetting rate. If $Y$ is a unit rate exponential random variable, then $\sigma$ is exponentially distributed with rate $r>0$.
The region covered by the resetting searcher up through time $t>0$ is
$S_{r}(t):=\cup_{s=0}^{t}B(X_{r}(s),R)\subseteq M,$
and the cover time of the nonempty compact target region $U_{T}\subseteq M$ is
$T_{r}:=\inf\{t>0:U_{T}\subseteq S_{r}(t)\}.$
We denote by $U_{T}^{\text{far}}\subseteq U_{T}$ the set of furthest target points, which means that
$L(x,X_{r}(0))>L(y,X_{r}(0))\quad\text{for all }x\in U_{T}^{\text{far}}\text{ and }y\in U_{T}\backslash U_{T}^{\text{far}}.$
Analogous to section 2 we define the first hitting time of $X_{r}(t)$ to a set $\Omega\subseteq M$,
$\tau_{r}(\Omega):=\inf\{t\geq 0:X_{r}(t)\in\Omega\}.$
We remark that throughout this work $T$ denotes a cover time and $\tau$ denotes a hitting time. If $\Omega\equiv B(x,R)$, then $\tau_{r}(\Omega)$ is equivalent to
$\tau_{r}(B(x,R))=\inf\{t>0:X_{r}(t)\in B(x,R)\}=\inf\{t>0:x\in S_{r}(t)\}.$
Finally, denote by $p_{\Omega}$ the probability of a stochastic searcher in $M\subseteq\mathbb{R}^{d}$ with initial position $X_{0}(0)=0$ hitting the set $\Omega\subseteq M$ before a resetting event at time $\sigma>0$,
$p_{\Omega}:=\mathbb{P}(\tau_{0}(\Omega)<\sigma).$
A key assumption in the results of this section is that for two sets $\Omega_{0},\Omega_{1}\subset M$,
(13) $L(\Omega_{0},X_{r}(0))>L(\Omega_{1},X_{r}(0))\quad\text{implies}\quad\lim_{r\to\infty}p_{\Omega_{0}}/p_{\Omega_{1}}=0$
where the distance between two sets $\Omega_{0},\Omega_{1}\subset M$ is given by
$L(\Omega_{0},\Omega_{1}):=\inf_{x\in\Omega_{0},y\in\Omega_{1}}L(x,y).$
In words, (13) means that a search is much more likely to hit a close target than a far target before a fast resetting time. This assumption is satisfied, for instance, by diffusive search with exponentially distributed resetting in which case $p_{\Omega}$ satisfies
$\ln p_{\Omega}\sim-\sqrt{L(\Omega_{0},X_{r}(0))^{2}r/D}\quad\text{as }r\to\infty$
where $D>0$ is the diffusivity (see Theorem 2 in [15]).
Now we state the main result of this section:
Theorem 1.
Under the assumptions of section 3, let $p^{*}=p^{*}(r)$ be any function of $r>0$ satisfying $p^{*}\sim p_{B(x^{*},R)}$ as $r\to\infty$ for any $x^{*}\in U_{T}^{\text{far}}$. Then
$\liminf_{r\to\infty}\mathbb{E}[(rp^{*}T_{r})^{m}]\geq m!\quad\text{as }r\to\infty$
for integers $m\geq 1$. Further,
$\limsup_{r\to\infty}\mathbb{E}[(rp_{B(x^{*},(1-\varepsilon)R)}T_{r})^{m}]=0\quad\text{for all }\varepsilon\in(0,1).$
In words, Theorem 1 says that in the frequent resetting limit we can bound a certain scaling of the cover time that depends only on the resetting rate and data on the set of furthest target points.
3.2. Proofs
Before we prove Theorem 1, we introduce a useful lemma which was recently proven (see Theorem 1 in [15]):
Lemma 2.
Under the assumptions of section 3, let $\Omega\subset M\backslash X_{r}(0)$ and let $p^{*}=p^{*}(r)$ be any function of $r>0$ satisfying $p^{*}\sim p_{\Omega}$ as $r\to\infty$. Then for integers $m\geq 1$, we have
$\mathbb{E}[(\tau_{r}(\Omega))^{m}]\sim\frac{m!}{(rp^{*})^{m}}\quad\text{as }r\to\infty.$
Proof of Theorem 1.
We start with the lower bound: Let $x^{*}\in U_{T}^{\text{far}}$. The survival probability of $T_{r}$ is that of the union of hitting times to each point in the target region, which itself has a natural lower bound,
(14) $\mathbb{P}(T_{r}>t)=\mathbb{P}(\cup_{x\in U_{T}}\{\tau_{r}(B(x,R))>t\})\geq\mathbb{P}(\tau_{r}(B(x^{*},R))>t).$
Hence, substituting $t$ for $z^{1/m}$ with $\0m\in\mathbb{N}\backslash\{0\}$ in (14) and integrating over $z\geq 0$ bounds the moments of $T_{r}$ from below,
$\mathbb{E}[(T_{r})^{m}]=\int_{0}^{\infty}$ $\mathbb{P}(T_{r}>z^{1/m})\,\textup{d}z$
$\geq\int_{0}^{\infty}\mathbb{P}(\tau_{r}(B(x^{*},R))>z^{1/m})\,\textup{d}z=\mathbb{E}[(\tau_{r}(B(x^{*},R)))^{m}].$
By Lemma 2 we find
$\liminf_{r\to\infty}\mathbb{E}[(rp_{B(x^{*},R)}T_{r})^{m}]\geq\lim_{r\to\infty}\mathbb{E}[(rp_{B(x^{*},R)}\tau_{r}(B(x^{*},R)))^{m}]=m!.$
Hence, for $p^{*}\sim p_{B(x^{*},R)}$ as $r\to\infty$, we have
$\liminf_{r\to\infty}\mathbb{E}[(rp^{*}T_{r})^{m}]\geq m!.$
Determining the limit supremum is more involved: Let $\varepsilon\in(0,1)$ and cover the compact target region $U_{T}\subseteq M$ using a finite set $\{x_{1},\dots,x_{k}\}\in M\backslash U_{T}^{\text{far}}$ such that the assumption in (13) is satisfied with $\Omega_{0}=U_{T}^{\text{far}}$ and $\Omega_{1}=\{x_{1},\dots,x_{k}\}$ and further that
$\cup_{i=1}^{k}B(x_{i},\varepsilon R)\supseteq U_{T}.$
With this cover, one can show [4] that for any $i\in\{1,\dots,k\}$,
$B(x_{i},\varepsilon R)\subset B(y_{i},R)\quad\text{if}\quad y_{i}\in B(x_{i},(1-\varepsilon)R).$
Hence,
(15) $\begin{split}\mathbb{P}(T_{r}>t)&=\mathbb{P}(\cup_{x\in U_{T}}\{\tau_{r}(B(x,R))>t\})\\ &\leq\mathbb{P}(\cup_{i=1}^{k}\{\tau_{r}(B(x_{i},(1-\varepsilon)R))>t\}).\end{split}$
Using the inclusion-exclusion principle [16] we rewrite the rightmost expression in (15),
(16) $\begin{split}\mathbb{P}(&\cup_{i=1}^{k}\{\tau_{r}(B(x_{i},(1-\varepsilon)R))>t\})\\ &=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{P}(\cap_{i\in I}\{\tau_{r}(B(x_{i},(1-\varepsilon)R))>t\})\Big{)}\\ &=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{P}(\min_{i\in I}\{\tau_{r}(B(x_{i},(1-\varepsilon)R))\}>t)\Big{)}.\end{split}$
Hence, substituting $t$ for $z^{1/m}$ with $\0m\in\mathbb{N}\backslash\{0\}$ in (15) and (16), and integrating over $z\geq 0$ bounds the moments of $T_{r}$ from above,
$\mathbb{E}[(T_{r})^{m}]$ $\leq\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\int_{0}^{\infty}\mathbb{P}(\min_{i\in I}\{\tau_{r}(B(x_{i},(1-\varepsilon)R))\}>z^{1/m})\,\textup{d}z\Big{)}$
$=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(\min_{i\in I}\{\tau_{r}(B(x_{i},(1-\varepsilon)R))\})^{m}]\Big{)}$
$=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}$
Multiplying through by $(rp_{B(x^{*},(1-\varepsilon)R)})^{m}$ yields
(17) $\begin{split}\mathbb{E}[(r&p_{B(x^{*},(1-\varepsilon)R)}T_{r})^{m}]\\ &\leq\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\mathbb{E}[(rp_{B(x^{*},(1-\varepsilon)R)}\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}\\ &=\sum_{j=1}^{k}\Big{(}(-1)^{j-1}\sum_{\begin{subarray}{c}I\subset\{1,\dots,k\}\\ |I|=j\end{subarray}}\Big{(}\frac{p_{B(x^{*},(1-\varepsilon)R)}}{p_{\cup_{i\in I}B(x_{i},(1-\varepsilon)R)}}\Big{)}^{m}\\ &\qquad\qquad\qquad\qquad\times\mathbb{E}[(rp_{\cup_{i\in I}B(x_{i},(1-\varepsilon)R)}\tau_{r}(\cup_{i\in I}B(x_{i},(1-\varepsilon)R)))^{m}]\Big{)}\end{split}$
Using Lemma 2 and the assumption in (13) to take the limit as $r\to\infty$ yields
(18) $\limsup_{r\to\infty}\mathbb{E}[(r$ $p_{B(x^{*},(1-\varepsilon)R)}T_{r})^{m}]\leq 0.$
Since the left-hand side of (18) is trivially non-negative, the proof is complete.
4. Cover times for a frequently resetting searcher on a network
4.1. Definitions and main result
Now consider a stochastic searcher on a discrete state space $\Gamma$ that randomly resets to its initial position. In particular, let $X=\{X(t)\}_{t\geq 0}$ be a non-resetting continuous-time Markov chain on a finite or countably infinite state space $\Gamma$. We encode the dynamics of $X$ in its infinitesimal generator matrix $Q=\{q(i,j)\}_{i,j\in\Gamma}$ whose off-diagonal entries give the (non-negative) rate that $X$ jumps from $i\in\Gamma$ to $j\in\Gamma$ [17]. By construction the diagonal entries of $Q$ are non-positive and are such that $Q$ has zero row sums. We assume that $\sup_{i\in\Gamma}|q(i,i)|<\infty$ so that $X$ cannot take infinitely many jumps in finite time.
Let $X_{r}$ be the process $X$ modified to instantaneously reset to its initial position $i_{0}\in\Gamma$ according to $\sigma$ as in (12). Assume the detection ‘radius’ of $X_{r}$ is zero; that is, the searcher sees only the node it occupies. Then the region covered by the resetting searcher up through time $t>0$ is
$S_{r}(t):=\cup_{s=0}^{t}X_{r}(s)\subseteq\Gamma,$
and the corresponding cover time of the target region $U_{T}\subseteq\Gamma$ is
$T_{r}:=\inf\{t>0:U_{T}\subseteq S_{r}(t)\}.$
As before, we define the first hitting time of $X_{r}(t)$ to a set $\Omega\subseteq\Gamma$,
$\tau_{r}(\Omega):=\inf\{t\geq 0:X_{r}(t)\in\Omega\},$
and the probability of a stochastic searcher hitting the set $\Omega\subseteq\Gamma$ before a resetting event at time $\sigma>0$,
$p_{\Omega}:=\mathbb{P}(\tau_{0}(\Omega)<\sigma).$
Following [18], define a path $\mathcal{P}$ of length $n\in\mathbb{N}$ from $i_{1}\in\Gamma$ to $i_{n+1}\in\Gamma$ to be a sequence of $n+1$ states in $\Gamma$,
$\mathcal{P}=(\mathcal{P}(1),\dots,\mathcal{P}(n+1))=(i_{1},\dots,i_{n+1})\in\Gamma^{n+1},$
so that
(19) $q(\mathcal{P}(j),\mathcal{P}(j+1))>0,\quad\text{for }j\in\{1,\dots,n\}.$
The condition in (19) says that there is a strictly positive probability of $X$ following the path $\mathcal{P}$. Naturally we assume there are paths from $i_{0}$ to each node in $U_{T}$.
For a path $\mathcal{P}\in\Gamma^{n+1}$, let $\lambda(\mathcal{P})$ be the product of the rates along the path,
(20) $\lambda(\mathcal{P}):=\prod_{i=1}^{n}q(\mathcal{P}(i),\mathcal{P}(i+1))>0.$
Let $d(\Gamma_{0},\Gamma_{1})\in\mathbb{N}$ denote the length of the geodesic path(s) from a set of nodes $\Gamma_{0}\subset\Gamma$ to a set of nodes $\Gamma_{1}\subset\Gamma$,
$d(\Gamma_{0},\Gamma_{1}):=\min\{n:\mathcal{P}\in\Gamma^{n+1},\mathcal{P}(1)\in\Gamma_{0},\mathcal{P}(n+1)\in\Gamma_{1}\}.$
We define the set of all paths from $\Gamma_{0}$ to $\Gamma_{1}$ with this minimum length $d(\Gamma_{0},\Gamma_{1})$,
$\mathcal{S}(\Gamma_{0},\Gamma_{1}):=\{\mathcal{P}\in\Gamma^{n+1}:\mathcal{P}(1)\in\Gamma_{0},\mathcal{P}(n+1)\in\Gamma_{1},n=d(\Gamma_{0},\Gamma_{1})\}.$
Define
(21) $\Lambda(\Gamma_{1}):=\sum_{\mathcal{P}\in\mathcal{S}(i_{0},\Gamma_{1})}\mathcal{P}(1)\lambda(\mathcal{P}).$
In words, the quantity $\Lambda(\Gamma_{1})$ in (21) is the sum of the products of the jump rates along the shortest paths from $i_{0}$ to $\Gamma_{1}$. If there is only one such path, then $\Lambda(\Gamma_{1})$ is merely the product of the jump rates along that path as in (20).
Finally, let $d^{*}\geq 1$ denote the length of the geodesic path(s) between $i_{0}$ and the furthest node(s) in the finite target set $U_{T}$,
$d^{*}:=\sup_{j\in U_{T}}d(i_{0},j),$
and let $U_{T}^{\text{far}}$ denote the furthest target point(s) in $U_{T}$,
$U_{T}^{\text{far}}:=\{j\in U_{T}:d(i_{0},j)=d^{*}\}\subseteq U_{T}.$
Unlike in previous sections, we do not require that there be a unique furthest target point in $U_{T}$ from the initial and resetting position, or equivalently that $|U_{T}^{\text{far}}|=1$.
With these definitions established, we now state the main result of this section:
Theorem 3.
Under the assumptions of section 4, for integers $m\geq 1$ we have
$\mathbb{E}[(T_{r})^{m}]\sim\frac{m!\,K_{m}}{r^{m(1-d^{*})}}\quad\text{as }r\to\infty$
where
$K_{m}:=\sum_{j=1}^{|U_{T}^{\text{far}}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T}^{\text{far}},\,|J|=j}(\Lambda(J))^{-m}\Big{)}.$
Theorem 3 above depends only on network properties along the shortest paths from the searcher initial position to the furthest parts of the target. If this furthest region contains only one node, we have the following corollary:
Corollary 4.
Under the assumptions of Theorem 3, assume further that $|U_{T}^{\text{far}}|=1$. Then
$\mathbb{E}[(T_{r})^{m}]\sim\frac{m!}{(rp_{0})^{m}}\quad\text{as }r\to\infty,$
where
$p_{0}=\frac{\Lambda(U_{T}^{\text{far}})}{r^{d^{*}}}\sim p_{U_{T}^{\text{far}}}\quad\text{as }r\to\infty.$
Further, $rp_{0}T_{r}$ converges in distribution to a unit rate exponential random variable,
$rp_{0}T_{r}\to_{\textup{dist}}\textup{Exponential}(1)\quad\text{as }r\to\infty.$
4.2. Proofs
Proof of Theorem 3.
By the inclusion-exclusion principle [16],
(22) $\begin{split}\mathbb{P}(T_{r}>t)&=\mathbb{P}(\cup_{j\in U_{T}}\{\tau_{r}(j)>t\})\\ &=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{P}(\cap_{j\in J}\{\tau_{r}(j)>t\})\Big{)}\\ &=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{P}(\tau_{r}(J)>t)\Big{)}\end{split}$
Substituting $t$ for $z^{1/m}$ with $\0m\in\mathbb{N}\backslash\{0\}$ in (22) and integrating over $z\geq 0$ gives an expression for the moments of $T_{r}$,
$\mathbb{E}[(T_{r})^{m}]=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\mathbb{E}[(\tau_{r}(J))^{m}]\Big{)}.$
For any set of nodes $J\in\Gamma$, it was recently shown that $p_{J}\sim\Lambda(J)/r^{d(i_{0},J)}$ as $r\to\infty$ [15]. This result combined with Lemma 2 and the assumption that $|U_{T}|<\infty$ yields
$\lim_{r\to\infty}\mathbb{E}[r^{m(1-d^{*})}T_{r}^{m}]$ $=\lim_{r\to\infty}\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\frac{r^{-md^{*}}}{p_{J}^{m}}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\lim_{r\to\infty}\frac{r^{-md^{*}}}{p_{J}^{m}}\lim_{r\to\infty}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=\sum_{j=1}^{|U_{T}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T},\,|J|=j}\lim_{r\to\infty}\frac{r^{m(d(i_{0},J)-d^{*})}}{(\Lambda(J))^{m}}$
×$\qquad\times\lim_{r\to\infty}\mathbb{E}[(rp_{J}\tau_{r}(J))^{m}]\Big{)}$
$=m!\sum_{j=1}^{|U_{T}^{\text{far}}|}\Big{(}(-1)^{j-1}\sum_{J\subseteq U_{T}^{\text{far}},\,|J|=j}\Lambda(J)^{-m}\Big{)}.$
The last step uses that $d(i_{0},J)<d^{*}$ unless $J\subseteq U_{T}^{\text{far}}$ in which case $d(i_{0},J)=d^{*}$. Setting $|U_{T}^{\text{far}}|=1$ immediately yields the moment convergence in Corollary 4 and convergence in distribution follows from Theorem 30.2 in [19]. ∎
5. Discussion
In this work we estimate all the moments of cover times of a wide range of stochastic search processes in $d$-dimensional continuous space and on an arbitrary discrete network in the frequent resetting limit. These results depend only on the resetting rate, which is well-defined for a large class of resetting time distributions, and a so-called success probability corresponding to the furthest target point(s). These probabilities need not be known exactly; their asymptotic behavior in the frequent resetting limit suffices. Recent work on success probabilities computes these values for a variety of stochastic search and resetting processes [20].
In the specific case of a one dimensional search under exponentially distributed resetting, we compute MCTs of an asymmetric interval for a Brownian search exactly and demonstrate their rapid convergence to the frequent resetting limit. This observation suggests that in reasonable domains the cover time of a resetting Brownian searcher with volume is closely approximated by our asymptotic results. On a symmetric interval, a prefactor of $3/2$ on the asymptotic result is explained by $p_{\cup_{i\in I}B(x_{i},(1-\varepsilon)R)}$ in (17) having frequent resetting behavior determined not merely by one endpoint of the target interval but by both ends. This result holds for both the unconstrained and constrained search.
Performing such exact computations in more complicated domains is involved or analytically intractable. Moreover, since cover times diverge in the frequent resetting limit, simulating this process is cumbersome. Hence the analytical results herein are all the more useful. That said, determining the role of dimension, domain geometry, and target geometry on these cover times would be a challenging and fruitful direction of future work.